# Expected Coordinate Improvement for High-Dimensional Bayesian Optimization

### v1

@@ 5 Conclusions
In this work, we propose a simple and efficient approach to extend the Bayesian optimization algorithm to high-dimensional problems. We propose the expected coordinate improvement criterion to measure the amount of improvement we can get by moving the current best solution along one coordinate. Based on the expected coordinate improvement, we propose a high-dimensional Bayesian optimization approach which optimizes one coordinate at each iteration following the order of the coordinates’ expected coordinate improvement values. The proposed algorithm shows very competitive performance on the selected test suite when compared with the standard Bayesian optimization algorithm and five state-of-the-art high-dimensional approaches. Applying the expected coordinate improvement approach to multiobjective optimization would be very interesting work for future research.

### v2

@@ 5 Conclusions
In this work, we propose a simple and efficient approach to extend the Bayesian optimization algorithm to high-dimensional problems. We propose the expected coordinate improvement (ECI) criterion to measure the amount of improvement when moving the current best solution along one coordinate. Based on the ECI, we propose a high-dimensional Bayesian optimization approach which optimizes one coordinate at each iteration following the order of the coordinates’ ECI values. The proposed algorithm shows very competitive performance on the selected test suite when compared with the standard Bayesian optimization algorithm and the state-of-the-art high-dimensional approaches. Our work provides an easy-to-understand and simple-to-implement approach for the Bayesian optimization field and has the potential to be used in solving real-world high-dimensional expensive optimization problems. However, our work does not consider expensive constraints and multiple expensive objectives. Applying the expected coordinate improvement approach to constrained optimization and multiobjective optimization would be very interesting work for future research.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62106207.